# Algorithmic contiguity from low-degree conjecture and applications in correlated random graphs

### v3

@@ Abstract
In this paper, assuming a natural strengthening of the low-degree conjecture, we provide evidence of computational hardness for two problems: (1) the (partial) matching recovery problem in the sparse correlated Erdős-Rényi graphs $\mathcal{G}(n,q;\rho)$ when the edge-density $q=n^{-1+o(1)}$ and the correlation $\rho<\sqrt{\alpha}$ lies below the Otter’s threshold, this resolves a remaining problem in [DDL23+]; (2) the detection problem between a pair of correlated sparse stochastic block model $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ and a pair of independent stochastic block models $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$ when $\epsilon^{2}\lambda s<1$ lies below the Kesten-Stigum (KS) threshold and $s<\sqrt{\alpha}$ lies below the Otter’s threshold, this resolves a remaining problem in [CDGL24+].
One of the main ingredient in our proof is to derive certain forms of algorithmic contiguity between two probability measures based on bounds on their low-degree advantage. To be more precise, consider the high-dimensional hypothesis testing problem between two probability measures $\mathbb{P}$ and $\mathbb{Q}$ based on the sample $\mathsf{Y}$. We show that if the low-degree advantage $\mathsf{Adv}_{\leq D}\big{(}\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=O(1)$, then (assuming the low-degree conjecture) there is no efficient algorithm $\mathcal{A}$ such that $\mathbb{Q}(\mathcal{A}(\mathsf{Y})=0)=1-o(1)$ and $\mathbb{P}(\mathcal{A}(\mathsf{Y})=1)=\Omega(1)$. This framework provides a useful tool for performing reductions between different inference tasks.
@@ 1 Introduction
Given an integer $n\geq 1$, for $(i,j)\in\operatorname{U}$ let $J_{i,j}$ and $K_{i,j}$ be independent Bernoulli variables with parameter $s$. In addition, let $\pi_{*}$ be an independent uniform permutation on $[n]=\{1,\dots,n\}$. Then, we define a triple of correlated random graphs $(G,A,B)$ as follows: first we generate $G$ independently with $\{I_{i,j},K_{i,j}\}$ and $\pi_{*}$ from a specific probability distribution over all graphs on $[n]$, and then (conditioned on $G$) we define for each $(i,j)\in\operatorname{U}$ that (note that we identify a graph with its adjacency matrix)
@@
Two fundamental problems in the study of correlated random graph model are as follows: (1) the detection problem, which involves determining whether a given pair of graphs $(A,B)$ is sampled from a pair of correlated random graphs or from a pair of independent random graphs; (2) the matching problem, which focuses on recovering the latent matching $\pi_{*}$ from a sample $(A,B)$ from the distribution of correlated random graphs. In recent years, significant progress has been made in understanding these problems for both the correlated Erdős-Rényi model and correlated stochastic block models. Through the collective efforts of the community, the information-theoretic thresholds for detection and matching have been fully characterized for correlated Erdős-Rényi model [CK17+, HM23, WXY22, WXY23, GML21, DD23a, DD23b, Du25+] and partially characterized for correlated SBMs [RS21, GRS22]. Additionally, various efficient detecting and matching algorithms have been developed with performance guarantees [DCKG19, BCL+19, DMWX21, FMWX23a, FMWX23b, GM20, GML24, GMS24, MRT21, MRT23, MWXY24, MWXY23, DL22+, DL23+, CR24, CDGL24+, Li24+]. Notably, like many other inference tasks in high-dimensional statistics [ZK16, RSS19, KWB22, Gam21], these problems appear to exhibit information-computation gaps. Specifically, for certain ranges of the correlation strength, detection or matching is information theoretically possible but no efficient algorithm is known to achieve these tasks. We now focus on the algorithmic side of these problems as they are more relevant to our work. Indeed, it has been shown that many inference tasks in the correlated random graph models exhibits sharp algorithmic phase transitions, as we elaborate below:
@@
  • For the detection problem between a pair of correlated Erdős-Rényi models $\mathcal{G}(n,q;\rho)$ and two independent Erdős-Rényi models $\mathcal{G}(n,q)$, we focus on the sparse regime where $q=n^{-1+o(1)}$. In this regime, on the one hand, it was shown in [MWXY24] that when $\rho>\sqrt{\alpha}$ where $\alpha\approx 0.338$ is the Otter’s constant [Ott48], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was shown in [DDL23+] that when $\rho<\sqrt{\alpha}$ there are evidence suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
@@
  • For the detection problem between a pair of correlated SBMs $\mathcal{S}(n,\frac{\lambda}{n};k,\epsilon;s)$ and two independent Erdős-Rényi graphs $\mathcal{G}(n,\frac{\lambda s}{n})$, we focus on the constant degree regime where $\lambda=O(1)$. In this regime, on the one hand, it was shown in [CDGL24+] that when $s>\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ where $\alpha\approx 0.338$ is the Otter’s constant and $\tfrac{1}{\epsilon^{2}\lambda}$ is the Kesten-Stigum threshold [KS66], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was also shown in [CDGL24+] that when $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ there are evidence suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
@@
    Can we provide the evidence that partial matching recovery is impossible (by efficient algorithms) in the same parameter regime where there are evidence suggesting that detection is impossible (by efficient algorithms)?
@@
We have the following results hold.
@@
    Assuming a natural strengthening of the low-degree conjecture (see Conjecture 2.5 for its precise meaning), then for the correlated Erdős-Rényi model $\mathcal{G}(n,q,\rho)$, when $q=n^{-1+o(1)}$ and $\rho<\sqrt{\alpha}$ it is impossible to recover a positive fraction of the coordinates of $\pi_{*}$ with constant probability by efficient algorithms.
@@
    Assuming a natural strengthening of the low-degree conjecture (see Conjecture 2.5 for its precise meaning), then for the correlated stochastic block models $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$, when $\lambda=O(1)$ and $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ it is impossible to recover a positive fraction of the coordinates of $\pi_{*}$ with probability tending to $1$ as $n\to\infty$ by efficient algorithms.
@@
    Assuming the low-degree conjecture (see Conjecture 2.2 for its precise meaning), then for the correlated stochastic block models $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$, when $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ it is impossible to strongly distinguish this model and a pair of independent SBMs $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$ by efficient algorithms, provided that the average degree $\lambda s$ is sufficiently large.
@@
Similarly, the results in [MWXY23, GM20, GML24, GMS24] naturally extends to show that there exists an efficient algorithm that achieves partial recovery of $\pi_{*}$ in a pair of correlated SBMs when $s>\sqrt{\alpha}$ with probability tending to $1$ as $n\to\infty$. Thus, Item (2) in Theorem 1.6 is also tight in the subcritical regime (i.e., when marginally two graphs are below the KS threshold).
@@
From Item (3) in Theorem 1.6 we see that for a pair of correlated SBMs $(A,B)$, when marginally both $A$ and $B$ are below the KS threshold (i.e., when $\epsilon^{2}\lambda s<1$), there is evidence suggesting that no efficient algorithm can strongly distinguish $(A,B)$ from a pair of independent stochastic block models when the correlation $s<\sqrt{\alpha}$. Since the result in [MWXY24] extends naturally to the case of stochastic block models which provides an efficient algorithm that strongly distinguish these two models when $s>\sqrt{\alpha}$, we see that in the subcritical regime (i.e., when marginally both $A$ and $B$ are below the KS threshold) the algorithmic correlation detection threshold is given by $\sqrt{\alpha}$. On the contrary, in the supercritical regime where $A$ and $B$ are above the KS threshold (i.e, when $\epsilon^{2}\lambda s>1$), we believe that the algorithmic correlation detection threshold should be strict lower than $\sqrt{\alpha}$.
@@ 1.1 Key challenges and innovations
    Define $g=g_{1}+\ldots+g_{n}$. We then have a statistics $g$ that can be efficiently computed, and (I) under $\mathbb{Q}$ we expect that $g$ is “large” as it is the sum of $n$ “of positive constant order” terms; (II) under $\mathbb{P}$ we expect that $g$ is “not large” as each $g_{i}$ should approximate $\mathbf{1}_{\{\pi_{*}(1)=i\}}$ in some sense. Thus, the statistics $g$ accumulates more signals than all low-degree polynomials, which (non-rigorously speaking) violates the low-degree conjecture.
@@
Another difficulty is that reducing between different inference tasks requires a framework that better constrains the behavior of efficiently computable estimators. To address this, we introduce a natural refinement of the low-degree conjecture: we posit that low-degree polynomials achieve signal-to-noise ratios that are at least as good as those of any efficient algorithm (see Conjecture 2.5 for a more precise description). While similar modifications to the low-degree conjecture have been explored in [MW23+], our motivations differ significantly. Their work focused on characterizing the limits of precise error rate for all computationally feasible algorithms, whereas our goal is to enable reductions between distinct inference tasks.
@@ 1.2 Comparison to concurrent work
In a concurrent work [DHSS25+], the authors investigate the computational hardness of the weak recovery problem in the stochastic block model $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon)$ when $k=n^{o(1)}$ and $\epsilon^{2}\lambda<1$ lies below the KS threshold. Notably, their proof is also based on a “recovery-to-detection reduction” approach, which argues that if there exists an efficient algorithm that achieves weak recovery, then we can use this algorithm to efficiently construct a statistics for detection (in a certain sense). Their results also relies on a strengthening of the low-degree conjecture, which is similar to the assumptions made in this paper.
However, the reduction technique employed in this paper differs significantly from that of [DHSS25+]. The authors of [DHSS25+] construct their statistic for detection using a correlation preserving projection technique established in [HS17]. This technique, which is designed specifically for weak recovery in block models, enables them to “regularize” the statistics and thus bound its moments under the null hypothesis. In contrast, our construction is more straightforward as we leverage the universal framework described in Section 1.1 to “transfer” the behavior of statistics across different probability measures. It seems that in comparison to [DHSS25+], our approach is easier to implement, less dependent on problem-specific details, and potentially applicable to a broader range of problems.
1.3 Notations
In this subsection, we record a list of notations that we shall use throughout the paper. Denote $\mathfrak{S}_{n}$ the set of permutations over $[n]$ and denote $\mu$ the uniform distribution on $\mathfrak{S}_{n}$. In addition, denote $\nu$ to be the uniform distribution on $[k]^{n}$. We will use the following notation conventions on graphs.
@@ 1.3 Notations
For two real numbers $a$ and $b$, we let $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$. We use standard asymptotic notations: for two sequences $a_{n}$ and $b_{n}$ of positive numbers, we write $a_{n}=O(b_{n})$, if $a_{n}<Cb_{n}$ for an absolute constant $C$ and for all $n$ (similarly we use the notation $O_{h}$ is the constant $C$ is not absolute but depends only on $h$); we write $a_{n}=\Omega(b_{n})$, if $b_{n}=O(a_{n})$; we write $a_{n}=\Theta(b_{n})$, if $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$; we write $a_{n}=o(b_{n})$ or $b_{n}=\omega(a_{n})$, if $a_{n}/b_{n}\to 0$ as $n\to\infty$. In addition, we write $a_{n}\circeq b_{n}$ if $a_{n}=[1+o(1)]b_{n}$. For a set $\mathsf{A}$, we will use both #$\#\mathsf{A}$ and $|\mathsf{A}|$ to denote its cardinality. For two probability measures $\mathbb{P}$ and $\mathbb{Q}$, we denote the total variation distance between them by $\operatorname{TV}(\mathbb{P},\mathbb{Q})$.

### v4

# Computational Lower Bounds for Correlated Random Graphs via Algorithmic Contiguity

@@ Abstract
In this paper, assuming the low-degree conjecture, we provide evidence of computational hardness for two problems: (1) the (partial) matching recovery problem in the sparse correlated Erdős-Rényi graphs $\mathcal{G}(n,q;\rho)$ when the edge-density $q=n^{-1+o(1)}$ and the correlation $\rho<\sqrt{\alpha}$ lies below the Otter’s threshold, this resolves a remaining problem in [DDL25]; (2) the detection problem between a pair of correlated sparse stochastic block models $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ and a pair of independent stochastic block models $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$ when $\epsilon^{2}\lambda s<1$ lies below the Kesten-Stigum (KS) threshold and $s<\sqrt{\alpha}$ lies below the Otter’s threshold, this resolves a remaining problem in [CDGL24+].
One of the main ingredient in our proof is to derive certain forms of algorithmic contiguity between two probability measures based on bounds on their low-degree advantage. To be more precise, consider the high-dimensional hypothesis testing problem between two probability measures $\mathbb{P}$ and $\mathbb{Q}$ based on the sample $\mathsf{Y}$. We show that if the low-degree advantage $\mathsf{Adv}_{\leq D}\big(\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big)=O(1)$, then (assuming the low-degree conjecture) there is no efficient algorithm $\mathcal{A}$ such that $\mathbb{Q}(\mathcal{A}(\mathsf{Y})=0)=1-o(1)$ and $\mathbb{P}(\mathcal{A}(\mathsf{Y})=1)=\Omega(1)$. This framework provides a useful tool for performing reductions between different inference tasks, without requiring a strengthened version of the low-degree conjecture as in [MW25, DHSS25].
@@ 1 Introduction
Given an integer $n\geq 1$, for $(i,j)\in\operatorname{U}$ let $J_{i,j}$ and $K_{i,j}$ be independent Bernoulli variables with parameter $s$. In addition, let $\pi_{*}$ be an independent uniform permutation on $n]=\{1,\dots,n\}$. Then, we define a triple of correlated random graphs $(G,A,B)$ as follows: first we generate $G$ independently with $\{J_{i,j},K_{i,j}\}$ and $\pi_{*}$ from a specific probability distribution over all graphs on $[n]$, and then (conditioned on $G$) we define for each $(i,j)\in\operatorname{U}$ that (note that we identify a graph with its adjacency matrix)
@@
Two fundamental problems in the study of correlated random graph model are as follows: (1) the detection problem, which involves determining whether a given pair of graphs $(A,B)$ is sampled from a pair of correlated random graphs or from a pair of independent random graphs; (2) the matching problem, which focuses on recovering the latent matching $\pi_{*}$ from a sample $(A,B)$ from the distribution of correlated random graphs. In recent years, significant progress has been made in understanding these problems for both the correlated Erdős-Rényi model and correlated stochastic block models. Through the collective efforts of the community, the information-theoretic thresholds for detection and matching have been fully characterized for correlated Erdős-Rényi model [CK17+, HM23, WXY22, WXY23, GML21, DD23a, DD23b, Du25+] and partially characterized for correlated SBMs [RS21, GRS22]. Additionally, various efficient detecting and matching algorithms have been developed with performance guarantees [DCKG19, BCL+19, DMWX21, FMWX23a, FMWX23b, GM20, GML24, GMS24, MRT21, MRT23, MWXY24, MWXY23, DL25, DL23+, CR24, CDGL24+, Li25]. Notably, like many other inference tasks in high-dimensional statistics [ZK16, RSS19, KWB22, Gam21], these problems appear to exhibit information-computation gaps. Specifically, for certain ranges of the correlation strength, detection or matching is information theoretically possible but no efficient algorithm is known to achieve these tasks. We now focus on the algorithmic side of these problems as they are more relevant to our work. Indeed, it has been shown that many inference tasks in the correlated random graph models exhibits sharp algorithmic phase transitions, as we elaborate below:
@@
  • For the detection problem between a pair of correlated Erdős-Rényi models $\mathcal{G}(n,q;\rho)$ and two independent Erdős-Rényi models $\mathcal{G}(n,q)$, we focus on the sparse regime where $q=n^{-1+o(1)}$. In this regime, on the one hand, it was shown in [MWXY24] that when $\rho>\sqrt{\alpha}$ where $\alpha\approx 0.338$ is the Otter’s constant [Ott48], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was shown in [DDL25] that when $\rho<\sqrt{\alpha}$ there is evidence suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
@@
  • For the detection problem between a pair of correlated SBMs $\mathcal{S}(n,\frac{\lambda}{n};k,\epsilon;s)$ and two independent Erdős-Rényi graphs $\mathcal{G}(n,\frac{\lambda s}{n})$, we focus on the constant degree regime where $\lambda=O(1)$. In this regime, on the one hand, it was shown in [CDGL24+] that when $s>\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ where $\alpha\approx 0.338$ is the Otter’s constant and $\tfrac{1}{\epsilon^{2}\lambda}$ is the Kesten-Stigum threshold [KS66], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was also shown in [CDGL24+] that when $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ there is evidence suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
@@
    Can we provide the evidence that partial matching recovery is impossible (by efficient algorithms) in the same parameter regime where there is evidence suggesting that detection is impossible (by efficient algorithms)?
@@
Assuming the low-degree conjecture (see Conjecture 2.2 for its precise meaning), we have the following results.
@@
    For the correlated Erdős-Rényi model $\mathcal{G}(n,q,\rho)$, when $q=n^{-1+o(1)}$ and $\rho<\sqrt{\alpha}$ it is impossible to recover a positive fraction of the coordinates of $\pi_{*}$ with constant probability by efficient algorithms.
@@
    For the correlated stochastic block models $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$, when $\lambda=O(1)$ and $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ it is impossible to recover a positive fraction of the coordinates of $\pi_{*}$ with probability tending to $1$ as $n\to\infty$ by efficient algorithms.
@@
    For the correlated stochastic block models $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$, when $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ it is impossible to strongly distinguish this model and a pair of independent SBMs $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$ by efficient algorithms, provided that the average degree $\lambda s$ is sufficiently large.
@@
Similarly, the results in [MWXY23, GM20, GML24, GMS24] naturally extend to show that there exists an efficient algorithm that achieves partial recovery of $\pi_{*}$ in a pair of correlated SBMs when $s>\sqrt{\alpha}$ with probability tending to $1$ as $n\to\infty$. Thus, Item (2) in Theorem 1.6 is also tight in the subcritical regime (i.e., when marginally two graphs are below the KS threshold).
@@
From Item (3) in Theorem 1.6 we see that for a pair of correlated SBMs $(A,B)$, when marginally both $A$ and $B$ are below the KS threshold (i.e., when $\epsilon^{2}\lambda s<1$), there is evidence suggesting that no efficient algorithm can strongly distinguish $(A,B)$ from a pair of independent stochastic block models when the correlation $s<\sqrt{\alpha}$. Since the result in [MWXY24] extends naturally to the case of stochastic block models which provides an efficient algorithm that strongly distinguish these two models when $s>\sqrt{\alpha}$, we see that in the subcritical regime (i.e., when marginally both $A$ and $B$ are below the KS threshold) the algorithmic correlation detection threshold is given by $\sqrt{\alpha}$. On the contrary, in the supercritical regime where $A$ and $B$ are above the KS threshold (i.e, when $\epsilon^{2}\lambda s>1$), we believe that the algorithmic correlation detection threshold should be strict lower than $\sqrt{\alpha}$. This was further supported by a recent work [CDGL25+], which has shown that in the case of $k=2$, strong detection between $\mathbb{P}_{n}$ and $\widetilde{\mathbb{Q}}_{n}$ is achievable when $\epsilon^{2}\lambda s>1$ and $s\geq\sqrt{\alpha}-\delta$ where $\delta$ is a sufficiently small constant. However, rigorous analysis in the general case seems of substantial challenge and we leave it for future work.
@@ 1.1 Key challenges and innovations
    Define $g=g_{1}+\ldots+g_{n}$. We then have a statistic $g$ that can be efficiently computed, and (I) under $\mathbb{Q}$ we expect that $g$ is “large” as it is the sum of $n$ “of positive constant order” terms; (II) under $\mathbb{P}$ we expect that $g$ is “not large” as each $g_{i}$ should approximate $\mathbf{1}_{\{\pi_{*}(1)=i\}}$ in some sense. Thus, the statistics $g$ should achieves detection in a certain sense, which violates our assumption.
@@ 1.2 Comparison to concurrent work
In a concurrent work [DHSS25], the authors investigate the computational hardness of the weak recovery problem in the stochastic block model $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon)$ when $k=n^{o(1)}$ and $\epsilon^{2}\lambda<1$ lies below the KS threshold. Notably, their proof is also based on a “recovery-to-detection reduction” approach, which argues that if there exists an efficient algorithm that achieves weak recovery, then we can use this algorithm to efficiently construct a statistic for detection (in a certain sense). Their results relies on a strengthening of the low-degree conjecture, which is inspired by the work [MW25].
However, the reduction technique employed in this paper differs significantly from that of [DHSS25]. The authors of [DHSS25] construct their statistic for detection using a correlation preserving projection technique established in [HS17]. This technique, which is designed specifically for weak recovery in block models, enables them to “regularize” the statistics and thus bound its moments under the null hypothesis. In contrast, our construction is more straightforward and only requires the original low-degree conjecture as we leverage the universal framework described in Section 1.1 to “transfer” the behavior of statistics across different probability measures. It seems that in comparison to [DHSS25], our approach is easier to implement, less dependent on problem-specific details, and potentially applicable to a broader range of problems.
1.3 Notation
In this subsection, we record a list of notations that we shall use throughout the paper. Denote $\mathfrak{S}_{n}$ the set of permutations over $[n]$ and denote $\mu$ the uniform distribution on $\mathfrak{S}_{n}$. In addition, denote $\nu$ to be the uniform distribution on $k]^{n}$. We will use the following notation conventions on graphs.
@@ 1.3 Notation
For two real numbers $a$ and $b$, we let $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$. We use standard asymptotic notations: for two sequences $a_{n}$ and $b_{n}$ of positive numbers, we write $a_{n}=O(b_{n})$, if $a_{n}<Cb_{n}$ for an absolute constant $C$ and for all $n$ (similarly we use the notation $O_{h}$ is the constant $C$ is not absolute but depends only on $h$); we write $a_{n}=\Omega(b_{n})$, if $b_{n}=O(a_{n})$; we write $a_{n}=\Theta(b_{n})$, if $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$; we write $a_{n}=o(b_{n})$ or $b_{n}=\omega(a_{n})$, if $a_{n}/b_{n}\to 0$ as $n\to\infty$. In addition, we write $a_{n}\circeq b_{n}$ if $a_{n}=[1+o(1)]b_{n}$. For a set $\mathsf{A}$, we will use both $\mathsf{A}$ and $|\mathsf{A}|$ to denote its cardinality. For two probability measures $\mathbb{P}$ and $\mathbb{Q}$, we denote the total variation distance between them by $\operatorname{TV}(\mathbb{P},\mathbb{Q})$.